# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

### v2

@@ Appendix C Global minima of representation cost for residual networks
$\mathscr{L}^{res}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y\\ \frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=1}^{L-2}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2},\quad\text{otherwise}.\end{cases}$
@@
Let $\sigma_{l}$ be a sequence of entrywise nonlinearities for which $\sigma_{l}(x)=0\iff x=0$ and assume intermediate widths satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$. Then, the global minima of the optimization problem
$\min_{Z^{1},\dots,Z^{L}}\mathscr{L}^{res}(Z)$
$\textnormal{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l}),\quad\forall\,l=1,\dots,L-2$
$\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,r_{L}$
@@
Let’s define with $\mathcal{S}$ the full constraint set and with $\mathcal{S}_{l}=\mathcal{S}_{l}(Z^{1},\dots,Z^{l-1},Z^{l+1},\dots,Z^{L})$ the constraint set on $Z^{l}$ with the other variables fixed. Notice that we can first minimize in $Z^{1}$ the problem
$\arg\min_{Z^{1}\in\mathcal{S}_{1}}\mathscr{L}^{res}(Z^{L},\dots,Z^{1})=\arg\min_{Z^{1}\in\mathcal{S}_{1}}\|\sigma_{2}^{-1}(Z^{2}-Z^{1})Z^{1,+}\|^{2}$
$\qquad\;\text{s.t.}\quad\sigma_{2}^{-1}(Z^{2}-Z^{1})Z^{1,+}Z^{1}=\sigma_{2}^{-1}(Z^{2}-Z^{1})$
whose minima is given by Lemma˜C.1 at $Z^{1,*}(Z^{2})=Z^{2}$ and minimal value equal to zero. Therefore, we can now minimize in $Z^{2}$ the representation cost
$\min_{Z^{2},\dots,Z^{L}}\mathscr{L}^{res}(Z^{L},\dots,Z^{2},Z^{1,*}(Z^{2}))=\frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=2}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2}$
$\text{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})\quad\forall\,l=2,\dots,L-1$
@@
Notice that the situation now is the same as the initial one, but with one layer less. Therefore, we can do the same thing for $Z^{2}$ using Lemma˜C.1 (with $g=\sigma_{l}^{-1}$) to obtain $Z^{2,*}(Z^{3})=Z^{3}$. By iterating this procedure for all layers, we arrive at
@@
$\text{s.t.}\quad Z^{L}=Y$
$\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
@@
(Stability of minima for ResNets) We highlight that the same $\Gamma$-convergence result as the one presented in Theorem˜4.4 can be stated also for ResNets, and the proof is exactly the same as the one presented in Section˜B.2. More precisely, we define
$\mathscr{L}_{\lambda}^{res}(Z)=\begin{cases}+\infty,\quad\text{if}\,\,\ Z^{L}\neq Y,\\ \frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=1}^{L-2}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2}+\frac{1}{\lambda}\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2},\quad\text{otherwise}.\end{cases}$ (27)
Then for $\lambda\to+\infty$, $\mathscr{L}_{\lambda}^{res}\to\mathscr{L}^{res}$ in the $\Gamma$-convergence sense, analogously to Theorem˜4.4. This will also be showcased numerically in Appendix˜D.
@@ Appendix D Additional numerical experiments
D.1 Additional values of $\delta$ for the singular value convergence
$\delta=10$ $\delta=8$
Refer to caption Refer to caption
$\delta=5$ $\delta=2$
Refer to caption Refer to caption
$\delta=1$ $\delta=10^{-1}$
Refer to caption Refer to caption
$\delta=10^{-2}$ $\delta=0$
Refer to caption Refer to caption
Figure 4: Singular values of trained networks against the predicted one from Lemmas˜B.1 and 4.3 for different values of $\delta=\frac{1}{\lambda}$ in Equation˜12. As we can observe, the singular values are converging to the predicted ones for $\delta\to 0$ as shown in Theorem˜$4.$
D.2 Residual networks
In this section we will show numerical results concerning Theorem˜C.2. The setting is the same of Section˜5, with $L=6$ layers (first and last feedforward, and the $4$ in the middle are residual). From Figure˜5, we can observe the evolution of the norm of the weights during optimization, also for varying $\delta=\frac{1}{\lambda}$ in Equation˜12. As we can see, for $\delta$ going to zero, all the weight matrices converge to zero as expected from Theorem˜C.2, but the first and the last ones, which are feedforward layers.
@@ D.2 Residual networks
Refer to caption Refer to caption
$\delta=2$ $\delta=1$
@@
Figure 5: Norm of the intermediate weights matrices for a for $L=6$ layer ResNet during training. As discussed in Remark˜C.3, we observe convergence to the structured minima discussed in Theorem˜C.2 for $\delta=\frac{1}{\lambda}$ (as in Equation˜12).
D.3 Visual representation of neural collapse
@@
Assume $Z$ satisfies NC1. For simplicity, we will denote with $z_{i}^{k}$ the $i$-th vector of the $k$-th class. Then we have
@@
$\alpha_{kk}=\langle z_{i}^{k},z_{j}^{k}\rangle,\quad\forall i,j=1,\dots,N_{k},\forall k=1,\dots,K.$
@@
$z_{i}^{k}=P^{\perp}_{\mathbb{A}^{k}}(0),\quad\forall i=1,\dots,N_{k},\forall k=1,\dots,K,$
where $\mathbb{A}_{k}=\text{Aff}(\{z_{j}^{k}\}_{j})$ is the affine subspace passing trhough the points of class $k$. By uniqueness of the orthogonal projection we get
$z_{i}^{k}=z_{j}^{k}=P_{\mathbb{A}^{k}}^{\perp}(0),\quad\forall i,j=1,\dots,N_{k},\forall k=1,\dots,K,$
@@
Given this premise, in the following experiments we will plot the Gram matrices $G_{l}=Z^{l,\top}Z^{l}$ for any $l=0,\dots,L-1$ in the case $L=15$ layer neural network trained in the setting of Section˜5. The emergence of the block structure is depicted in Figure˜6.
Moreover, we sampled a random plane in each feature space, i.e., for any $l$ we sample a random orthogonal matrix $B^{l}\in\mathbb{R}^{2\times n_{l}},B^{l}B^{l,\top}=I_{2}$. Figure˜7 shows the projection of the feature matrices $Z^{l}$ on the plane spanned by the rows of $B^{l}$, i.e. $B^{l}Z^{l}$, for which we plot all the columns as points in $\mathbb{R}^{2}$, allowing us to pictorially see the deep-neural (rank) collapse in all intermediate layers in this random subspace.

### v3

@@ Appendix C Global Minima of Representation Cost for Residual Networks
$\mathscr{L}^{res}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y\\ \frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=2}^{L-2}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2},\quad\text{otherwise}.\end{cases}$
@@
Let $\sigma_{l}$ be a sequence of entrywise nonlinearities for which $\sigma_{l}(x)=0\iff x=0$ and assume intermediate widths satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$. Then, for almost every $Z^{1}=\sigma_{1}(W_{1}X)$, the global minima of the optimization problem
$\min_{Z^{2},\dots,Z^{L}}\mathscr{L}^{res}(Z)$
$\quad\textnormal{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l}),\quad\forall\,l=1,\dots,L-2$
$\quad\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\quad\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,r_{L}$
@@
Let’s define with $\mathcal{S}$ the full constraint set and with $\mathcal{S}_{l}=\mathcal{S}_{l}(Z^{1},\dots,Z^{l-1},Z^{l+1},\dots,Z^{L})$ the constraint set on $Z^{l}$ with the other variables fixed. Notice that we can first minimize in $Z^{2}$ the problem
$\arg\min_{Z^{2}\in\mathcal{S}_{2}}\mathscr{L}^{res}(Z^{L},\dots,Z^{1})=\arg\min_{Z^{2}\in\mathcal{S}_{2}}\|\sigma_{3}^{-1}(Z^{3}-Z^{2})Z^{2,+}\|^{2}$
$\qquad\;\text{s.t.}\quad\sigma_{3}^{-1}(Z^{3}-Z^{2})Z^{2,+}Z^{2}=\sigma_{3}^{-1}(Z^{3}-Z^{1})$
whose minima is given by Lemma˜C.1 at $Z^{2,*}(Z^{3})=Z^{3}$ and minimal value equal to zero. Therefore, we can now minimize in $Z^{3}$ the representation cost
$\min_{Z^{3},\dots,Z^{L}}\mathscr{L}^{res}(Z^{L},\dots,Z^{2},Z^{2,*}(Z^{3}))=\frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=3}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2}$
$\text{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})\quad\forall\,l=3,\dots,L-1$
@@
Notice that the situation now is the same as the initial one, but with one layer less. Therefore, we can do the same thing for $Z^{3}$ using Lemma˜C.1 (with $g=\sigma_{l}^{-1}$) to obtain $Z^{3,*}(Z^{4})=Z^{4}$. By iterating this procedure for all layers, we arrive at
@@
$\quad\text{s.t.}\quad Z^{L}=Y$
$\quad\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\quad\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,K.$
@@
(Stability of minima for ResNets) We highlight that the same $\Gamma$-convergence result as the one presented in Theorem˜4.6 can be stated also for ResNets, and the proof is exactly the same as the one presented in Section˜B.5. More precisely, given $\phi:\mathcal{S}\to\mathbb{R}_{+}$ lower semicontinuous, we define
[…]
Then for $\lambda\to+\infty$, $\mathscr{L}_{\lambda}^{res}\to\mathscr{L}^{res}$ in the $\Gamma$-convergence sense, analogously to Theorem˜4.6. This will also be showcased numerically in Appendix˜D.
@@ Appendix D Additional Numerical Experiments
D.1 Residual Networks
In this section we will show numerical results concerning Theorem˜C.2. The setting is the same of Section˜5, with $L=6$ layers (first and last feedforward, and the $4$ in the middle are residual). From Figure˜4, we can observe the evolution of the norm of the weights during optimization, also for varying $\delta=\frac{1}{\lambda}$ in Equation˜13. As we can see, for $\delta$ going to zero, all the weight matrices converge to zero as expected from Theorem˜C.2, but the first and the last ones, which are feedforward layers.
@@ D.1 Residual Networks
Figure 4: Norm of the intermediate weights matrices for a for $L=6$ layer ResNet during training. As discussed in Remark˜C.3, we observe convergence to the structured minima discussed in Theorem˜C.2 for $\delta=\frac{1}{\lambda}$ (as in Equation˜13).
D.2 Visual Representation of Neural Collapse
@@
For simplicity, throughout the proof, we will denote with $z_{i}^{k}$ the $i$-th vector of the $k$-th class. Assume first that $Z$ satisfies NC1. Then we have that $z_{i}^{k}=z^{k}$ for all $i=1,\dots,N_{k}$ and
@@
$\alpha_{kk}=\langle z_{i}^{k},z_{j}^{k}\rangle,\quad\forall i,j=1,\dots,N_{k}\;\text{and}\quad\forall k=1,\dots,K.$
@@
$z_{i}^{k}=P^{\perp}_{\mathbb{A}^{k}}(0),\quad\forall i=1,\dots,N_{k}\;\text{and}\quad\forall k=1,\dots,K,$
where $\mathbb{A}_{k}=\text{Aff}(\{z_{j}^{k}\}_{j})$ is the affine subspace passing through the points of class $k$. By uniqueness of the orthogonal projection we get
$z_{i}^{k}=z_{j}^{k}=P_{\mathbb{A}^{k}}^{\perp}(0)\quad\forall i,j=1,\dots,N_{k}\;\text{and}\quad\forall k=1,\dots,K,$
@@
Given this premise, in the following experiments we will plot the Gram matrices $G_{l}=Z^{l,\top}Z^{l}$ for any $l=0,\dots,L-1$ in the case of a $L=15$ layer neural network trained in the setting of Section˜5. The emergence of the block structure is depicted in Figure˜5.
Moreover, we sampled a random plane in each feature space, i.e., for any $l$ we sample a random orthogonal matrix $B^{l}\in\mathbb{R}^{2\times n_{l}},B^{l}B^{l,\top}=I_{2}$. Figure˜6 shows the projection of the feature matrices $Z^{l}$ on the plane spanned by the rows of $B^{l}$, i.e. $B^{l}Z^{l}$, for which we plot all the columns as points in $\mathbb{R}^{2}$, allowing us to pictorially see the deep-neural (rank) collapse in all intermediate layers in this random subspace.